4. Assume the following information:

Work-in process, beginning 100 units, 25% complete

Units completed and transferred out 600 units

Work-in-process, ending 200 units, 40% complete

A. Compute the equivalent units of production using the weighted-average method.

B. Compute the equivalent units of production using the FIFO method.
Answer #1
A
Weighted-average method:
Units completed and transferred out 600
Work-in-process, ending 80 =200*40%
Equivalent units of production 680
B
FIFO method:
Work-in process, beginning 75 =100*(1-25%)
Units started and completed 500 =600-100
Work-in-process, ending 80
Equivalent units of production 655